OpenStudy (anonymous):

Could this be modified to shift to the left/right on a graph? $\left\{\left|x\right|<2:\sqrt{\cos x}\cos \left(300x\right)-.1+\sqrt{\left|x\right|}\right\}$

OpenStudy (turingtest):

\[\left\{\left|x\right|<2:\sqrt{\cos x}\cos \left(300x\right)-.1+\sqrt{\left|x\right|}\right\}\]?

OpenStudy (anonymous):

Yes. It's an equation I found, when you graph it it makes a heart. I wanted to make a design on Desmos that involves having two connected hearts, but I don't know how you could left/right shift that.

OpenStudy (anonymous):

https://www.desmos.com/ If seeing what it is helps ^-^

OpenStudy (turingtest):

\[\left\{\left|x\right|<2:\sqrt{\cos x}\cos \left(300x\right)-0.1+\sqrt{\left|x\right|}\right\}\]I am interpreting as the function and interval\[f(x)=\sqrt{\cos x}\cos \left(300x\right)-0.1+\sqrt{|x|},\{x:-2\le x\le2\}\]

OpenStudy (turingtest):

I don't know much what to make of that link, but if you ignore the bounds, then putting\(f(x-h)\) will shift the graph to the right bu h units

OpenStudy (anonymous):

Alright, I'll try that ^u^

OpenStudy (anonymous):

Yay! It works!!! Thx :D